If a stack of 20 is 1/8 inch, how much would a stack of 750 be?
Answer by nerdybill(7384)

If a stack of 20 is 1/8 inch, how much would a stack of 750 be?
.
(1/8)/20 = x/750
cross multiplying:
20x = (1/8)(750)
x = (1/8)(750)(1/20)
x = (1/8)(75)(1/2)
x = 75/16
x = 4 and 11/16 inches